Required information The Foundational 15 (Algo) (L04-1, LO4-2, LO4-3, LO4-4, LO4-5) (The following information applies to the5. Compute the equivalent units of production for conversion. Equivalent units of production6. What is the cost of beginning work in process inventory plus the cost added during the period for materials? Total cost of7. What is the cost of beginning work in process inventory plus the cost added during the period for conversion? Total cost o8. What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.) Cost per equivalent unit for9. What is the cost per equivalent unit for conversion? (Round your answer to 2 decimal places.) Cost per equivalent unit10. What is the cost of ending work in process inventory for materials? (Round your intermediate calculations to 2 places.) C
Answer #1
5.) Conversion
Units transferred out (4,700 + 37,200 - 8,200 )         33,700
Add: Work in process , June 30 ( 50% ) - 8,200           4,100
Equivalent units of production         37,800
6.) Total cost of materials $ 161,315 =16040+145275
7.) Total cost of conversion $ 215,460 =12960+92500+110000
8.) Materials
Units transferred out (4,700 + 37,200 - 8,200 )         33,700
Add: Work in process , June 30 ( 100% ) - 8,200           8,200
Equivalent units of production         41,900
Cost per equivalent unit for materials $ 3.85 =161315/41900
9.) Cost per equivalent unit for Conversion $ 5.70 =215460/37800
10.) Cost of ending work in process inventory for materials $ 129,745 =33700*3.85